OpenStudy (angiepangie0726):

The function f(x) = 4(3)x represents the growth of a dragonfly population every year in a remote swamp. Erin wants to manipulate the formula to an equivalent form that calculates four times a year, not just once a year. Which function is correct for Erin's purpose, and what is the new growth rate?

OpenStudy (angiepangie0726):

i have the answer choices too if you think thatll help??

OpenStudy (angiepangie0726):

i have the answer choices too if you think thatll help??

OpenStudy (irishboy123):

if you have the answers to hand, this is the clue: \(\large f(x) = 4(3)^x = 4(3)^\frac{4x}{4}\) they want you to re-write the question so that you are compounding a quarterly rate 4 times over the period of a year. use \((a^n)^m = a^{n \times m}\) i think the question is really badly written, and really quite pointless to, so you have my sympathy,.... for what it's worth :p